OpenStudy (anonymous):

f(x) = 8-3lx+2l find f(-6) ? Show how to get the answer.

OpenStudy (anonymous):

plug -6 for x |-6+2| = 4 because of the absolute value 8-3(4) -3*4= -12 8-12= -4

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

no problem! find someone to help me with my problem its literally impossible :(